Consider the following demand for a certain product:

Period

Demand

1

25

2

120

3

40

4

60

5

30

6

140

7

60

8

80

9

35

10

150

11

55

12

90

  1. Develop a forecast for these data, using an exponential smoothing model? What is the resulting MSD?
  2. Develop a forecast for these data, using an exponential smoothing with a linear trend model? What is the resulting MSD?
  3. Develop a forecast for these data, using Winters method? What is the resulting MSD?
  4. Which method from the methods mentioned above is “better”? How do you know that?
Answer #1

1. exponential smoothing model

Period D=Demand F=Forecast Deviation Square of Devaiation
1 25
2 120 25.00 95.00 9025.00
3 40 44.00 -4.00 16.00
4 60 43.20 16.80 282.24
5 30 46.56 -16.56 274.23
6 140 43.25 96.75 9360.95
7 60 62.60 -2.60 6.75
8 80 62.08 17.92 321.17
9 35 65.66 -30.66 940.22
10 150 59.53 90.47 8184.75
11 55 77.62 -22.62 511.86
12 90 73.10 16.90 285.63
MSD 2655.35
F+1 = (Dxα)+((1-α)F)
α = 0.2
1-α= 0.8

2. exponential smoothing with a linear trend

Period D=Demand L=Level T=Trend F=Forecast Deviation Square of Devaiation
0 54.55 2.95
1 25 51.00 1.00 57.50 32.50 1056.25
2 120 65.60 5.08 52.00 -68.00 4623.38
3 40 64.55 3.24 70.69 30.69 941.75
4 60 66.23 2.78 67.79 7.79 60.74
5 30 61.21 0.43 69.01 39.01 1521.78
6 140 77.31 5.14 61.64 -78.36 6139.84
7 60 77.96 3.79 82.45 22.45 504.02
8 80 81.40 3.68 81.75 1.75 3.06
9 35 75.07 0.68 85.08 50.08 2508.40
10 150 90.60 5.13 75.75 -74.25 5513.61
11 55 87.59 2.69 95.73 40.73 1659.05
12 90 90.22 2.67 90.28 0.28 0.08
889.68 MSD 2134.16
F+1 = (Dxα)+((1-α)F)
α = 0.2
1-α= 0.8
β = 0.3
1-β = 0.7
Coefficients
Intercept 54.5454545
X Variable 1 2.95454545

through regression analysis

3. Winters Method

Period D=Demand Deseasonalized Demand Deseasonalized Demand
(After Regression)
Seasonlized Factor
(No initialization)
Seasonlized Factor L=Level T=Trend F=Forecast Deviation Square of Devaiation
0 54.88 2.94
1 25 57.82 0.43 0.43 57.85 2.95 24.94 -0.06 0.00
2 120 60.76 1.98 1.90 61.30 3.10 115.22 -4.78 22.83
3 40 61.875 63.70 0.63 0.68 63.21 2.74 44.09 4.09 16.71
4 60 65 66.64 0.90 0.97 65.09 2.48 64.17 4.17 17.40
5 30 70 69.58 0.43 0.43 67.96 2.60 29.17 -0.83 0.68
6 140 75 72.51 1.93 1.93 70.98 2.73 135.91 -4.09 16.70
7 60 78.125 75.45 0.80 0.66 77.18 3.77 48.55 -11.45 131.08
8 80 80 78.39 1.02 0.95 81.65 3.98 76.69 -3.31 10.93
9 35 80.625 81.33 0.43 0.44 84.53 3.65 37.38 2.38 5.68
10 150 81.25 84.27 1.78 1.95 85.94 2.98 171.90 21.90 479.79
11 55 87.21 0.63 0.72 86.45 2.24 63.84 8.84 78.22
12 90 90.15 1.00 0.96 89.63 2.52 85.46 -4.54 20.60
897.34 MSD 72.78
F+1 = (Dxα)+((1-α)F)
α = 0.2
1-α= 0.8
β = 0.3
1-β = 0.7
ϒ = 0.5
p= 4
Coefficients
Intercept 54.88095
X Variable 1 2.938988

through regression analysis

Conclusion: Out of all 3 methods, winters method is more suitable because MSD is least of all 3 .i.e 72.78 that means it has less variation & more accurate forecasting.

(Alpha, Beta & Gamma as well as P are assumed)